OpenStudy (anonymous):

-9e+3.1+5e-7.1=8 what does e equal

OpenStudy (anonymous):

first combine like terms, get all the numbers to one side and all the variables to one side. -9e + 5e = -4e 3.1 - 7.1 = -4 so you have -4e - 4.1 = 8 add 4.1 to each side -4e = 12.1 divide both sides by -4 e = -12.1/4

OpenStudy (anonymous):

k thanks